OpenStudy (anonymous):

Solving Quadratic Equations: what is the square root of -7x²=-448 ? show work please :/

OpenStudy (anonymous):

there is nothing to work out, the index \[\sqrt[n]{x}\]n being the index, is even and the radicand, x, is negative, so that is not a real number

OpenStudy (anonymous):

well what about taking the square root of x²=-2x+33 can that be worked out?

OpenStudy (anonymous):

that would be \[x=\sqrt{-2x+33}\]

OpenStudy (chris):

hmmm, how about solving algebraically. So 1. -7x^2 = -448 2. (-7x^2)/-7 = -448/-7 //divide each side by -7 3. x^2 = 64 4. Squrt(x^2) = Squrt(64) //take the square root of both sides to get rid of the exponent 5. x = 8

OpenStudy (chris):

or does it have to be done as a quadratic?

OpenStudy (anonymous):

hmm.. my assignment says solving quadratic functions and to take the square root of the number I listed above.

OpenStudy (anonymous):

@Confucious where do I go from there?

OpenStudy (chris):

hmm, I think in the case of solving as a quadratic: 1. -7x^2 = -448 2. -7x + 0x +448 = 0 3. So in the form of ax^2 + bx +c = 0, a = -7, b = 0, c = 448.

OpenStudy (anonymous):

oh that's it, you leave it like that

OpenStudy (chris):

heh also I was wrong btw. x^2 = 64 means x=8 or x=-8

OpenStudy (chris):

so now plug in a=-7, b = 0, and c=448 into http://en.wikipedia.org/wiki/Quadratic_equation - both variations (plus and minus), to solve for x

OpenStudy (mr.math):

For your first problem, do as chris said \(-7x²=-448 \implies x^2=64 \implies x=\pm 8\).

OpenStudy (anonymous):

are you doing nonreal numbers? do you know about \[i\]?

OpenStudy (mr.math):

For the second problem \(x²=-2x+33 \iff x^2+2x-33=0\), you need to use the quadratic formula \(\large x={-b \pm \sqrt{b^2-4ac} \over 2a}\), with a=1, b=2 and c=-33.

OpenStudy (anonymous):

@Chris you are sort of confusing me.. I got this part: 1. -7x^2 = -448 2. -7x + 0x +448 = 0 3. So in the form of ax^2 + bx +c = 0, a = -7, b = 0, c = 448. but not your second set of directions?

OpenStudy (mr.math):

@Confucious: The problem has solutions in real numbers.

OpenStudy (anonymous):

wait my answer was wrong, use the quadratic formula to solve it, you can do it that way

OpenStudy (chris):

Like Mr. Math said (if I'm reading the question correctly) - if your goal is to solve for x, the form of ax^2 + bx + c = 0 allows you to use the quadratic equation Mr. Math posted to solve for both solutions of x

OpenStudy (chris):

just by plugging in the numbers a, b, and c

OpenStudy (anonymous):

@Confucious so that's it? You don't take the square root because you can't? I'm unsure about the unreal numbers.

OpenStudy (anonymous):

@Mandy17, use the quadratic formula like Mr. Math said and you can solve it

OpenStudy (anonymous):

@Mr. Math, which problem are you talking about?

OpenStudy (chris):

mandy - read the first two sections of this page - http://en.wikipedia.org/wiki/Quadratic_equation Basically if you can write a formula in the form of : ax^2 + bx + c = 0, where x is unknown and a, b, and c are known, you can use the quadratic formula to solve for x - the unknown.

OpenStudy (chris):

the idea is, given an equation - use algebra to put zero on one side of it, and see if you have something that resembles ax^2+bx + c

OpenStudy (anonymous):

and my answer I come up with from the quadratic formula is my answer to: taking the square root?

OpenStudy (chris):

in the first problem, we did: 1. -7x²=-448 2. -7x^2 +448 = -448 + 448 //add 448 to both sides so that zero is on the right hand side 3. -7x^2 + 448 = 0 4. Now look to see if you have something in the form of ax^2 + bx + c = 0. You do! 5. a = -7 (look above), since there is nothing that's just multiplied by x ("bx"), you still have a "b", but it's zero, and c is 448. 6. i.e. -7x^2 + 0x + 448 = 0, and since this is the format of a quadratic equation, we can use a = -7, b = 0, c = 448 , and stick them into the formula that Mr. Math put in.

OpenStudy (chris):

you should get -8 and +8 (since really the quadratic formula always has you do two cases - "b +/-" means "b plus or minus" - so you have to plug in the numbers twice. One where you add to b, one where you subtract.

OpenStudy (anonymous):

@Christ the final total outcome of the quadratic formula is -8 and +8? or is there more I need to do.